# Origin of unexpected weak Gilbert damping in the LSMO/Pt bilayer system

### v3

@@ Abstract
We present a first-principles and semiclassical analysis of the recent puzzling observation that a La0.7Sr0.3MnO3 (LSMO) thin film displays larger Gilbert damping than an LSMO/Pt bilayer, the opposite of what conventional spin-pumping theory predicts. Density-functional calculations combined with Wannier interpolation give an intrinsic damping of $\alpha_{\mathrm{int}}^{\mathrm{LSMO}}\!\approx\!1.4\times 10^{-3}$, confirming that the experimentally reported high damping must be extrinsic. Guided by the self-induced inverse spin Hall effect (ISHE) already demonstrated in LSMO by Gupta et al. [Phys. Rev. B 109, 014437 (2024)], we quantify how the large spin Hall angle $|\theta_{\mathrm{SH}}|\simeq 0.093$ and the small longitudinal conductivity of LSMO lead to efficient spin-to-charge conversion inside the ferromagnet, enhancing the effective damping. For the LSMO/Pt heterostructure, we find that Pt capping strongly increases the longitudinal conductivity, reducing the interfacial $|\theta_{\mathrm{SH}}|$ to 0.007 and thus suppressing spin-pumping-induced damping. By applying a Valet–Fert analysis to our layer-resolved ab-initio spin accumulation, we extract the spin-diffusion length within Pt, which, together with the calculated conductivities and spin Hall angle, yields a non-negligible anti-damping spin–orbit torque coefficient. This coefficient indicates the potential for an antidamping spin-orbit torque (SOT) to reduce the net damping in current-biased LSMO/Pt. The observed reduction in effective damping of the LSMO/Pt bilayer thus arises from the competition between the intrinsic Gilbert damping of the LSMO film and an anti-damping torque that becomes prominent when spin pumping is suppressed by the small interfacial spin Hall angle. Our results reconcile the apparently anomalous damping hierarchy without invoking any new interfacial mechanisms, and provide a quantitative design rule via the tunable ratio of Hall to longitudinal conductivities and spin-diffusion length for engineering magnetically robust, low-loss manganite heterostructures for spin-orbitronic applications.
@@ I Introduction
Our study demonstrates that the experimental observation of smaller damping in the Pt/LSMO system compared to pure LSMO can be explained by the significant reduction in the spin Hall angle ($\theta_{SH}$). We show that in the Pt/LSMO system, the presence of the metallic Pt layer increases the longitudinal conductivity ($\sigma_{xx}$), resulting in a much lower $\theta_{SH}$. This substantial reduction in $\theta_{SH}$ decreases both the spin-orbit torque (SOT) antidamping effects and the damping effect due to the spin-pumping effect. However, the balance between the conversion from charge to spin (anti-damping) and spin to charge (damping) in this heterostructure leads to an overall smaller damping.
@@ III.3 Longitudinal charge ($\sigma_{xx}$) and spin Hall ($\sigma_{SHC}$) conductivities
The longitudinal conductivity ($\sigma_{xx}$) and spin Hall conductivity ($\sigma_{xy}$) for both systems are depicted in Figure 2. The $\sigma_{xx}$ plots (Figures 2a and 2c) highlight a stark contrast between the two systems. For LSMO, the small longitudinal conductivity is observed at zero chemical potential ($\mu=0$), indicating limited conducting channels. In contrast, the presence of Pt in LSMO/Pt significantly enhances $\sigma_{xx}$ at zero chemical potential due to the additional pathways for charge transport which can be further understood from the Fig.1 where it is clear that Pt introduces additional states near Fermi energy facilitating better charge transport. As can be seen from the figure, the longitudinal conductivity ($\sigma_{xx}$) for LSMO is very small at $\mu=0$, indicating poor charge transport properties near the Fermi level ($E_{F}$). This suggests a a low density of states (DOS) at the Fermi level which is consistent with the electronic band structure shown in Fig.1 which depicts a single band crossing the Fermi level. However, it is important to note that at higher chemical potentials, LSMO has larger $\sigma_{xx}$.
It may be noted here that, although the interface between LSMO and Pt may exhibit finite roughness and a 1.3% lattice mismatch, these effects are not expected to overshadow the pronounced disparity in longitudinal conductivities. The additional scattering arising from interface imperfections could modify the transport behavior, but the substantial conductivity difference of over an order of magnitude suggests that our main conclusions remain vaild.
@@
[…]
@@ III.4 Self-induced ISHE in LSMO and its impact on effective damping
The observed Gilbert damping in LSMO is larger than in LSMO/Pt, which can be rationalized by considering the details of the spin-orbit torque mechanisms. It is important to note that recently Gupta et al. reported observation of a self-induced inverse spin Hall effect (ISHE) and spin pumping [48] within LSMO. It can be noted that similar self-induced ISE was observed by Tsukahara et al. [49] in permalloy at room temperature. This self-induced ISHE leads to the generation of a significant spin-pumping voltage within the LSMO film, indicating the presence of appreciable spin-to-charge conversion within the material. In the case of LSMO the absence of an external interface means there is no anti-damping mechanism from an external source (due to the injection of spin current from another material). The self-induced spin pumping effect powered by larger values of $|\theta_{\text{SH}}|$ in LSMO, therefore, leads to significant angular momentum dissipation within the material itself, contributing to larger effective damping via conversion of spin current to charge current through $\mathbf{J}_{c}=\theta_{SH}\left(\frac{\hbar}{2e}\right)(\mathbf{J}_{s}\times\mathbf{\sigma})$. Here $\mathbf{J}_{c}$ and $\mathbf{J}_{s}$ are respectively charge and spin current density. $\mathbf{\sigma}$ is the direction of the spin polarization [50].
Another important point we can gather from the sign of the spin Hall angle we have computed. The sign of $\theta_{SH}$ determines the direction of the charge current generated by the ISHE. It can be seen that $\theta_{SH}$ for LSMO is negative. However, the absolute value $|\theta_{SH}|$ indicates the strength of the spin-to-charge conversion. Thus, a large $|\theta_{SH}|$ (regardless of whether $\theta_{SH}$ is positive or negative) can result in a significant ISHE. Recently, Bleser et al. [51] demonstrated large spin-to-charge conversion with a negative spin Hall angle in thermally evaporated chromium thin films. The effective sign of $\theta_{SH}$ can also be influenced by the thickness of the film. The difference in the composition between the MnO2 and (La,Sr)O layers in the structural model we considered (see Fig.1) therefore induces a strong SOC at the interface. This gradient can enhance the ISHE in a manner similar to FePt films as seen in recent experiments [52].
@@ III.5 Spin-diffusion length in Pt and antidamping SOT in LSMO/Pt
In the linear-response regime, where the induced spin accumulation is small, it is reasonable to assume that the DFT-calculated spin polarization $P(z)$ is proportional to $\mu_{s}(z)$. The general solution to Eq. (7) for a film of thickness $t_{\text{Pt}}$, with appropriate boundary conditions (e.g., $z=0$ at the LSMO/Pt interface and $z=t_{\text{Pt}}$ at the Pt/vacuum interface), can be expressed in a hyperbolic cosine form [57, 58],
@@
In the experimental setup by Gupta et al. [7], the Pt layer thickness was 3 nm. A direct first-principles treatment of such a thick slab would be computationally prohibitive. Therefore, our DFT calculations employed a thinner 4-monolayer Pt film ($t_{\text{Pt}}\approx 0.8$ nm). Initially, an exponential decay model was considered for fitting the DFT-derived $P(z)$ data, which yielded $\lambda_{\text{sf}}\approx 2.175$ nm. However, this result ($\lambda_{\text{sf}}\approx 2.7\times t_{\text{Pt}}$) indicated that the condition $t_{\text{Pt}}\gg\lambda_{\text{sf}}$ was not met, making the exponential approximation inappropriate for our DFT slab. Therefore, we again fitted the layer-resolved DFT polarization data from the 4 ML Pt slab directly to the more general hyperbolic cosine model given by Eq. (8). This fitting procedure yielded a spin-diffusion length of $\lambda_{\text{sf,Pt}}\approx 0.7975$ nm ($7.975$). This value, where $\lambda_{\text{sf,Pt}}\approx t_{\text{Pt}}$, confirms the necessity of using the hyperbolic model for our computationally constrained thin Pt slab. While this $\lambda_{\text{sf,Pt}}$ is shorter than some bulk Pt values, it is a plausible result for an ultrathin film where interface scattering and confinement effects can significantly influence spin transport.
@@
×$\quad\times\frac{|\tilde{G}^{\uparrow\downarrow}|^{2}+\text{Re}[\tilde{G}^{\uparrow\downarrow}]\tanh^{2}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}{|\tilde{G}^{\uparrow\downarrow}|^{2}+2\text{Re}[\tilde{G}^{\uparrow\downarrow}]\tanh^{2}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})+\tanh^{4}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}$ (9)
[…]
@@ Acknowledgments
This work was supported by the Korea Institute of Science and Technology, GKP (Global Knowledge Platform, Grant number 2V6760) project of the Ministry of Science, ICT and Future Planning.

### v4

@@ Abstract
This study presents a first-principles and semiclassical analysis of the puzzling observation that a La0.7Sr0.3MnO3 (LSMO) thin film exhibits larger Gilbert damping than an LSMO/Pt bilayer, contrary to conventional spin-pumping expectations. Density-functional theory with Wannier interpolation yields an intrinsic damping of $\alpha_{\mathrm{int}}^{\mathrm{LSMO}}\!\approx\!1.4\times 10^{-3}$, supporting an extrinsic origin of the high experimental value. Guided by the self-induced inverse spin Hall effect (ISHE) demonstrated in LSMO [Gupta et al., Phys. Rev. B 109, 014437 (2024)], we argue that the large spin Hall angle $|\theta_{\mathrm{SH}}|\simeq 0.093$ and low longitudinal conductivity of LSMO enable an efficient conversion of spin current to charge current boosting the effective damping. In the LSMO/Pt heterostructure the Pt cap shunts the charge current, raising $\sigma_{xx}$ and reducing the interfacial $|\theta_{\mathrm{SH}}|$ to 0.007. A Valet–Fert analysis for layer-resolved ab-initio spin accumulation gives the Pt spin-diffusion length and a non-negligible antidamping SOT coefficient, qualitatively accounting for the observed damping reduction under current bias. The seemingly anomalous damping hierarchy is thus reconciled without invoking additional interfacial mechanisms. The distinct length scales governing spin-pumping normalization, namely, the short absorption depth relevant to self-pumping in a single LSMO film versus the full magnetic thickness applicable to an LSMO/Pt bilayer, are crucial in this context. This observation suggests a practical design strategy: by simultaneously tuning the spin Hall-to-longitudinal conductivity ratio and the spin-diffusion length, one can engineer heterostructures with minimized magnetic losses for spin-orbitronics applications.
@@ I Introduction
Our study proposes the experimental observation of higher damping pure LSMO can be related to the higher $\theta_{SH}$ in the LSMO films compared to the Pt/LSMO system where the presence of the metallic Pt layer increases the longitudinal conductivity ($\sigma_{xx}$), resulting in a much lower $\theta_{SH}$.
@@ III.2 Intrinsic damping
It is important to acknowledge that while LSMO demonstrates half-metallicity as temperature approaches zero, its spin polarization is observed to decrease due to factors such as structural imperfections, non-zero temperatures, and defects introduced during growth. Point-contact Andreev reflection and spin-resolved photoemission studies generally report spin polarization values of approximately $P\simeq 0.75{-}0.85$ at 300 K for epitaxial films on $\mathrm{SrTiO}_{3}$. In contrast, polycrystalline samples may exhibit values below 0.6. [42, 43]
@@ III.3 Longitudinal charge ($\sigma_{xx}$) and spin Hall ($\sigma_{SHC}$) conductivities
The longitudinal conductivity ($\sigma_{xx}$) and spin Hall conductivity ($\sigma_{xy}$) for both systems are depicted in Figure 2. The $\sigma_{xx}$ plots (Figures 2a and 2c) highlight a stark contrast between the two systems. For LSMO, the small longitudinal conductivity is observed at zero chemical potential ($\mu=0$), indicating limited conducting channels. In contrast, the presence of Pt in LSMO/Pt significantly enhances $\sigma_{xx}$ at zero chemical potential due to the additional pathways for charge transport which can be further understood from the Fig.1 where it is clear that Pt introduces additional states near the Fermi energy facilitating better charge transport. As can be seen from the figure, the longitudinal conductivity ($\sigma_{xx}$) for LSMO is very small at $\mu=0$, indicating poor charge transport properties near the Fermi level ($E_{F}$). This suggests a low density of states (DOS) at the Fermi level which is consistent with the electronic band structure shown in Fig.1 which depicts a single band crossing the Fermi level. However, it is important to note that at higher chemical potentials, LSMO has larger $\sigma_{xx}$.
It may be noted here that, although the interface between LSMO and Pt may exhibit finite roughness and a 1.3% lattice mismatch, these effects are not expected to overshadow the pronounced disparity in longitudinal conductivities. The additional scattering arising from interface imperfections could modify the transport behavior, but the substantial conductivity difference of over an order of magnitude suggests that our main conclusions remain valid.
@@
This indicates that the relaxation time in LSMO is approximately 2.1 times longer than in LSMO/Pt. Furthermore, we recalculated the longitudinal charge conductivity for LSMO using a relaxation time of $\tau=21$ fs, which resulted in a conductivity of $27738.05\,\text{S/m}$, compared to $13869.02\,\text{S/m}$ calculated earlier with $\tau=10$ fs. Despite this increase, LSMO’s conductivity remains significantly lower than that of LSMO/Pt, which is $295584.22\,\text{S/m}$—approximately 11 times larger.
@@ III.4 Self-induced ISHE in LSMO and its impact on effective damping
The observed Gilbert damping in LSMO is larger than in LSMO/Pt, which can be rationalized by considering the details of the spin-orbit torque mechanisms. It is important to note that recently Gupta et al. reported observation of a self-induced inverse spin Hall effect (ISHE) and spin pumping [50] within LSMO. It can be noted that similar self-induced ISHE was observed by Tsukahara et al. [51] in permalloy at room temperature. This self-induced ISHE leads to the generation of a significant spin-pumping voltage within the LSMO film, indicating the presence of appreciable spin-to-charge conversion within the material. In the case of LSMO the absence of an external interface means there is no anti-damping mechanism from an external source (due to the injection of spin current from another material). The self-induced spin pumping effect powered by larger values of $|\theta_{\text{SH}}|$ in LSMO, therefore, leads to significant angular momentum dissipation within the material itself, contributing to larger effective damping via conversion of spin current to charge current through $\mathbf{J}_{c}=\theta_{SH}\left(\frac{\hbar}{2e}\right)(\mathbf{J}_{s}\times\mathbf{\sigma})$. Here $\mathbf{J}_{c}$ and $\mathbf{J}_{s}$ are charge and spin current density, respectively. $\mathbf{\sigma}$ is the direction of the spin polarization [52].
Another important point we can gather from the sign of the spin Hall angle we have computed: the sign of $\theta_{SH}$ determines the direction of the charge current generated by the ISHE. It can be seen that $\theta_{SH}$ for LSMO is negative. However, the absolute value $|\theta_{SH}|$ indicates the strength of the spin-to-charge conversion. Thus, a large $|\theta_{SH}|$ (regardless of whether $\theta_{SH}$ is positive or negative) can result in a significant ISHE. Recently, Bleser et al. [53] demonstrated large spin-to-charge conversion with a negative spin Hall angle in thermally evaporated chromium thin films. The effective sign of $\theta_{SH}$ can also be influenced by the thickness of the film. The difference in the composition between the MnO2 and (La,Sr)O layers in the structural model we considered (see Fig.1) therefore induces a strong SOC at the interface. This gradient can enhance the ISHE in a manner similar to FePt films as seen in recent experiments [54].
@@ III.5 Spin-diffusion length in Pt and antidamping SOT in LSMO/Pt
In the linear-response regime, where the induced spin accumulation is small, it is reasonable to assume that the DFT-calculated spin polarization $P(z)$ is proportional to $\mu_{s}(z)$ with slowly varying profiles along z.. The general solution to Eq. (7) for a film of thickness $t_{\text{Pt}}$, with appropriate boundary conditions (e.g., $z=0$ at the LSMO/Pt interface and $z=t_{\text{Pt}}$ at the Pt/vacuum interface), can be expressed in a hyperbolic cosine form [59, 60],
@@
In the experimental setup by Gupta et al. [7], the Pt layer thickness was 3 nm. A direct first-principles treatment of such a thick slab would be computationally prohibitive. Therefore, our DFT calculations employed a thinner 4-monolayer Pt film ($t_{\text{Pt}}\approx 0.8$ nm). Initially, an exponential decay model was considered for fitting the DFT-derived $P(z)$ data, which yielded $\lambda_{\text{sf}}\approx 2.175$ nm. However, this result ($\lambda_{\text{sf}}\approx 2.7\times t_{\text{Pt}}$) indicated that the condition $t_{\text{Pt}}\gg\lambda_{\text{sf}}$ was not met, making the exponential approximation inappropriate for our DFT slab. Therefore, we again fitted the layer-resolved DFT polarization data from the 4 ML Pt slab directly to the more general hyperbolic cosine model given by Eq. (8). This fitting procedure yielded a spin-diffusion length of $\lambda_{\text{sf,Pt}}\approx 0.7975$ nm ($7.975$ Å). This value, where $\lambda_{\text{sf,Pt}}\approx t_{\text{Pt}}$, confirms the necessity of using the hyperbolic model for our computationally constrained thin Pt slab. While this $\lambda_{\text{sf,Pt}}$ is shorter than some bulk Pt values, it is a plausible result for an ultrathin film where interface scattering and confinement effects can significantly influence spin transport.
@@
$\quad\times\frac{|\tilde{G}^{\uparrow\downarrow}|^{2}+\text{Re}[\tilde{G}^{\uparrow\downarrow}]\tanh^{2}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}{|\tilde{G}^{\uparrow\downarrow}|^{2}+2\text{Re}[\tilde{G}^{\uparrow\downarrow}]\tanh^{2}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})+\tanh^{4}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}$ (9)
Where the dimensionless scaled spin-mixing conductance, $\tilde{G}^{\uparrow\downarrow}$, is defined through $\tilde{G}^{\uparrow\downarrow}=G^{\uparrow\downarrow}_{\text{unscaled}}\frac{2\lambda_{\text{sf,Pt}}\tanh(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}{\sigma_{\text{eff}}}$. Using Pt thickness, $t_{\text{Pt}}=$0.8\text{\,}\mathrm{n}\mathrm{m}$\$ , spin-diffusion length in Pt, $\lambda_{\text{sf,Pt}}\approx$0.8\text{\,}\mathrm{n}\mathrm{m}$\$ (from our DFT-based hyperbolic fit as explained above), spin Hall angle of Pt, $|\theta_{\text{SH}}|=0.007$ (obtained from our DFT calculations for the LSMO/Pt system), Interfacial spin-mixing conductance parameter, $g^{\uparrow\downarrow}=1.488\times 10^{19}$ $\mathrm{m}^{-2}$ (adapted from the experimental work of Gupta et al. [7] on a similar LSMO/Pt bilayer) we obtained the value of $\tilde{G}^{\uparrow\downarrow}$. This value is converted to the unscaled conductance $G^{\uparrow\downarrow}_{\text{unscaled,Re}}=g^{\uparrow\downarrow}(e^{2}/h)\approx 5.765\times 10^{14}\,$\mathrm{\SIUnitSymbolOhm}^{-1}\,\mathrm{m}^{-2}$\$. The imaginary part, $\text{Im}[G^{\uparrow\downarrow}_{\text{unscaled}}]$, is assumed to be zero for this estimation. To calculate the effective conductivity in the equation for $\tilde{G}^{\uparrow\downarrow}$ for the $\sigma_{\text{eff}}$, we utilize the total longitudinal conductivity of our LSMO(2ML)/Pt(4ML) bilayer obtained from our Boltzmann transport calculations, which is $\sigma_{\text{bilayer}}=$295\,584.22\text{\,}\mathrm{S}\mathrm{/}\mathrm{m}$\$. With these factors, we obtain the magnitude of the antidamping torque coefficient to be estimated as $|\tau_{ad}|\approx 0.002$.
@@ Acknowledgments
This work was supported by the Korea Institute of Science and Technology, GKP (Global Knowledge Platform, Grant number 2V6760) project of the Ministry of Science, ICT and Future Planning. S.B and P.G.thank the Department of Atomic Energy (DAE) , Govt. of India.
Appendix A Spin-pumping and spin-orbit torque contributions to the effective Gilbert damping
The effective Gilbert damping in the presence of both spin-pumping and spin-orbit torque (SOT)-induced anti-damping is given by
$\alpha_{\mathrm{eff}}=\alpha_{\mathrm{int}}+\Delta\alpha_{\mathrm{sp}}-\Delta\alpha_{\mathrm{ad}}.$ (10)
For the bare LSMO film, the pumped spins are reabsorbed inside LSMO within one spin-diffusion length, whereas for the bilayer they leave the ferromagnet and are absorbed in Pt. Consequently, the spin-pumping mechanisms in the two cases differ.
For the LSMO/Pt interface, spin pumping follows the conventional description [63]
$\Delta\alpha_{\mathrm{sp}}^{\text{LSMO/Pt}}=\frac{\gamma\hbar}{4\pi M_{s}\,t_{\text{LSMO}}}\,g^{\uparrow\downarrow}_{\!\mathrm{eff}},$ (11)
where the heavy metal (Pt) acts as a spin sink. The spin-pumping-induced damping $\Delta\alpha_{\mathrm{sp}}$ depends weakly on the spin Hall angle $\theta_{\mathrm{SH}}$ through spin-backflow correction factor, $\mathcal{R}_{\mathrm{bf}}$,
$\Delta\alpha_{\mathrm{sp}}=\frac{\gamma\hbar}{4\pi M_{s}t_{\mathrm{FM}}}g^{\uparrow\downarrow}_{\mathrm{eff}},\quad g^{\uparrow\downarrow}_{\mathrm{eff}}=g^{\uparrow\downarrow}(1-\mathcal{R}_{\mathrm{bf}}),\quad\mathcal{R}_{\mathrm{bf}}\propto\theta_{\mathrm{SH}}^{2}.$ (12)
For a bare LSMO film, there is no spin sink; the angular momentum must be absorbed within the material. Following Gupta et al. [50], $\alpha_{\mathrm{eff}}=\alpha_{\mathrm{int}}+\alpha_{\mathrm{sp}}\propto\sqrt{\theta_{\mathrm{SH}}}$, which can be approximated as
$\alpha\approx\sqrt{\frac{\theta_{\mathrm{SH}}\,w(2e/\hbar)\lambda_{LSMO}\tanh\!\left(\frac{t}{2\lambda_{LSMO}}\right)}{I_{\mathrm{ISHE}}}\frac{g_{r}^{\uparrow,\downarrow}\gamma B_{\mathrm{rf}}^{2}\hbar}{4\pi\mu_{0}M_{s}}}.$ (13)
Here $I_{\mathrm{ISHE}}$ is the measured inverse spin Hall current and $\lambda_{LSMO}$ is the spin-diffusion length in the LSMO film.
While the above approach is more physically motivated, it requires accurate transferability between computationally obtained parameters and experimental conditions. Moreover, certain parameters like $B_{rf}$ (the rf magnetic field associated with microwaves) are experimental in nature and challenging to assume from the theory.
A simpler alternative is to use Eq. (11) with a different length scale. In the self-pumping case, the angular momentum is absorbed within the ferromagnet, and we may introduce an effective absorption depth $\lambda_{\mathrm{abs}}$ equal to the measured spin-diffusion length in LSMO ($\approx 0.8$ nm) determined from hot-electron attenuation spectroscopy [64]:
$\Delta\alpha_{\mathrm{sp}}^{\text{LSMO}}=\frac{\gamma\hbar}{4\pi M_{s}\,\lambda_{\text{abs}}}\,g^{\uparrow\downarrow}_{\!\mathrm{eff}}.$ (14)
To calculate the anti-damping contribution numerically, we may start with the interfacial spin–orbit torque density localized at the FM||HM interface (z=0z=0[61],
$\mathbf{T}(z)=\delta(z)\,\frac{g\mu_{B}\,J_{c}}{2e}\left[\tau_{ad}\,\hat{\mathbf{m}}\times(\hat{\mathbf{m}}\times\hat{\mathbf{y}})+\tau_{f}\,\hat{\mathbf{m}}\times\hat{\mathbf{y}}\right$], (15)
where $J_{c}$ is the in-layer charge current density in the heavy metal, and $\tau_{{ad},f}$ are the dimensionless (anti)damping-like and field-like torque efficiencies, respectively. For a ferromagnet of thickness $t_{F}$, the spatially averaged torque density is $\overline{\mathbf{T}}=\mathbf{T}_{\rm int}/t_{F}$. Balancing this against the LLG torque $\gamma\mu_{0}M_{s}\,\hat{\mathbf{m}}\times(\hat{\mathbf{m}}\times\mathbf{H}_{AD})$ and using $g\mu_{B}=\hbar\gamma$ gives the effective anti-damping (damping-like) field
$\boxed{\;\mu_{0}H_{AD}\;=\;\frac{\hbar}{2e}\,\frac{J_{c}\,\tau_{d}}{M_{s}\,t_{F}}\;}.$ (16)
The damping-like efficiency we have already mentioned in the main text ($\tau_{ad}=0.002$).
To express this as a scalar change in the Gilbert damping, we map the anti-damping field to a linewidth-equivalent increment under small-angle steady precession at angular frequency $\omega$,
$\boxed{\Delta\alpha_{ad}\;\approx\;\frac{\gamma\,\mu_{0}H_{AD}}{\omega}\,\cos\varphi\;},$ (17)
where $\varphi$ is the phase between the SOT drive and $\dot{\hat{\mathbf{m}}}$ set by the microwave circuit and device geometry. In the absence of an independent phase calibration, taking $\varphi=0$ (i.e. $\cos\varphi=1$) provides an upper-bound estimate. For typical measurements one may use $f=\omega/2\pi=10~\text{GHz}$.
Using realistic experimental parameters, the calculated contributions to the total damping are summarized in Table 1. While the theoretical values do not exactly match the experiment, they reproduce the qualitative trends.
Table 1: Calculated contributions to the damping and comparison with experiment (f=10 GHz, $J_{c}=10^{11}$ A/m2).
Sample $\theta_{\mathrm{SH}}$ $\alpha_{\mathrm{int}}$ $\Delta\alpha_{\mathrm{sp}}$ $\Delta\alpha_{\mathrm{ad}}$ $\alpha_{\mathrm{eff}}$ (calc./exp.)
LSMO 0.093 0.0014 0.0788 0 0.0802 / 0.0104
LSMO/Pt 0.007 0.012 0.0064 0.005 0.0134 / 0.0046
In the calculations, we used the experimental $g^{\uparrow\downarrow}$, with $\theta_{\mathrm{SH}}$, $M_{s}$, and $\lambda_{\mathrm{sf}}$ obtained from our ab initio calculations.